OpenStudy (anonymous):

find (complex) zeros f(x)=x^4+5x^2+4

OpenStudy (anonymous):

f(x)=x^4+5x^2+4 = (x^2 + 4) (x^2 + 1) x = +/-2i, +/-i

OpenStudy (solomonzelman):

it's easier than you think.

OpenStudy (solomonzelman):

\[f(x)=x^4+5x^2+4\]\[Let~~x^2=a\]\[f(x)=a^2+5a+4\]\[a^2+5a+4=0\]\[(a+1)(a+4)=0\]\[a=-1~~~,~~~~a=-4\]\[x^2=-1~~~~,~~~x^2=-4\]\[x=±~i~~~~,~~~x=±~2i\]

OpenStudy (solomonzelman):

See how I'm doing it, no knowledge besides imaginary numbers and basic algebra.